OpenStudy (anonymous):

Need Help PLease?? The population in a small city is expanding and its current population (P) is given by the equation P = P0(1.13))^n where P0 is 125,000 and n is 14. What is the city's current population? A. 691,844 B. 227,500 C. 1,977,500 D. 492

OpenStudy (texaschic101):

P = 125,000(1.13^14) P = 125,000(5.5348) P = 691,850 --- its a little off because I rounded, but your answer is A